OpenStudy (anonymous):

Need the steps to solve this inequality:

OpenStudy (anonymous):

\[x^4-x \le 0\]

OpenStudy (michele_laino):

hint: we have to factorize the binomial at the left side

OpenStudy (michele_laino):

for example, at the first step, we can write: \[\Large x\left( {{x^3} - 1} \right) \leqslant 0\]

OpenStudy (michele_laino):

now, we have to factorize x^3-1, do you know how to factorize it?

OpenStudy (anonymous):

Yes

OpenStudy (michele_laino):

ok! Then rewrite my inequality above, using your factorization

OpenStudy (anonymous):

would it be (x-1)( x+1)( x -1)?

OpenStudy (michele_laino):

not exactly, we have this: \[\Large x\left( {x - 1} \right)\left( {{x^2} + x + 1} \right) \leqslant 0\]

OpenStudy (michele_laino):

since: \[\Large {x^3} - 1 = \left( {x - 1} \right)\left( {{x^2} + x + 1} \right)\]

OpenStudy (anonymous):

I see

OpenStudy (michele_laino):

now, we can note that \[\Large {{x^2} + x + 1}\] is always positive

OpenStudy (michele_laino):

so we have to study the sign of x and x-1 only

OpenStudy (michele_laino):

for example, please solve this inequality: \[\Large x - 1 \geqslant 0\]

OpenStudy (anonymous):

\[x \ge 1\]

OpenStudy (michele_laino):

correct! So we have this drawing: 

OpenStudy (michele_laino):

a continuous line stands for positivity, whereas a dashed line stands for negativity

OpenStudy (michele_laino):

so we have the subsequent drawing: